4. 013 points I Previous Answers LarLinAlg8 2.5.009 Use the matrix of transition probabilities P and initial state matrix Xo to find the state matrices X1, 2, and X3. T 0.6 0.2 0.1 T 0.1 P 0.3 0.7 0.1 Xo 0.2 0.1 0.1 0.8
Answer #1

When we multiply the initial state matrix X0 with the transition matrix P, we get state matrix X1.
X_{1}=PX_{0}

Again multiplying with the transition matrix, we get state matrix X2 and so on.
X_{2}=PX_{1}
X_{3}=PX_{2}

\therefore X_{1}=\begin{bmatrix} 0.6 & 0.2 & 0.1\\ 0.3& 0.7 &0.1 \\ 0.1& 0.1 & 0.8 \end{bmatrix}\begin{bmatrix} 0.1\\ 0.2\\ 0.7 \end{bmatrix}= \begin{bmatrix} 0.17\\ 0.24\\ 0.59 \end{bmatrix}

X_{2}=\begin{bmatrix} 0.6 & 0.2 & 0.1\\ 0.3& 0.7 &0.1 \\ 0.1& 0.1 & 0.8 \end{bmatrix}\begin{bmatrix} 0.17\\ 0.24\\ 0.59 \end{bmatrix}= \begin{bmatrix} 0.209\\ 0.278\\ 0.513 \end{bmatrix}